OpenStudy (anonymous):

i need an algebraic equation with only one variable for the following word problem: A jar contains $18.50 in dimes and quarters. There are 110 coins in the jar. How many dimes and quarters are there?

OpenStudy (ashleyisakitty):

where x=number of quarters, and 110-x=number of dimes (because the number of dimes is 110 minus the number of quarters): .25(x) + .10(110-x) = 18.50 .25x + 11 - .10x = 18.50 .25x - .10x = 7.50 .15x = 7.50 x = 50 50 quarters, 60 dimes

OpenStudy (anonymous):

Oh my. That makes sense, thank you (: